40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Message ID: 1206     Entry time: Fri Feb 20 05:41:04 2015
Author: Zach 
Type: DailyProgress 
Category: Laser 
Subject: Simple ISS test 

I wanted to do some intensity feedback testing, for two reasons:

  1. Just to get used to using the fiber amplitude modulators
  2. While I wait for the machined parts for the 1064nm M2 ISS testing on the old gyro table, I might as well use this basically perfect setup to do some initial runs with the M2 at 1550nm

So that I can have as much power as possible, I removed the fiber phase modulator and installed the amplitude modulator in its place. To generate the PDH sidebands, I simply drove into the laser bias tee with the 30 MHz oscillator signal and increased the amplitude until I got the same modulation depth as I measured with the modulator. I also had to readjust the demod phase via cable lengths, but after that the cavity locked just as before (and with an identical OLTF---not shown here). I don't claim that this locking technique is as good as using a phase modulator, in light of possible RFAM effects, but it is likely fine for intensity testing.

I also tried to increase the DC drive current of the laser, but it kept stalling after I tried to increase it above ~115 mA (the output power would increase in accordance with the plot on the datasheet, but then would suddenly crash and not return if the current was lowered until the driver output ON/OFF was cycled---not sure what gives here). So, I set it to 100 mA, where it seemed stable. The output of the laser head at this current is ~12 mW, so the max-transmission output of the amplitude modulator is about 6 mW (due to the 50% insertion loss). Adding a slight DC offset to the modulator, I reduced the output to ~92% to get some linear actuation strength for feedback.

I then tried to create an AC-coupled loop with an SR560, but had problems with stability on the low end. Eventually, I gave up and used the A-B function to subtract the measured DC level of around 4 V from the TRANS PD signal. I then put a pole at 300 Hz and scaled up the gain until I saw oscillations up near 100 kHz, and then slightly back down. Using this offset-subtracted DC-coupled loop, I was able to get solid in-loop performance, obtaining a UGF near 100 kHz and suppressing fluctuations to the dark noise level (consistent with the PDA255's noise) over a wide band.

The next step will be to use my low-noise readout optoelectronics and try out the Chachi servo.

 

ELOG V3.1.3-